OpenStudy (anonymous):

To which numbers could you apply cancelling when calculating the product 8/9 * 5/6?

OpenStudy (anonymous):

\[\frac{8}{9} * \frac{5}{6}\] 2 can be removed from 8 and 6 \[=\frac{4}{9} *\frac{5}{3}\]